OpenStudy (anonymous):

the length of a rectangular floor is 2 feet more than its width . the area of the floor is 168 square feet . kim wants to use a rug in the middle of the room and leave a 2-foot border of the floor visible on all sides .what should the width (the shorter side ) of the rug be ?

OpenStudy (anonymous):

the length of a rectangular floor is 2 feet more than its width . the area of the floor is 168 square feet . kim wants to use a rug in the middle of the room and leave a 2-foot border of the floor visible on all sides .what should the width (the shorter side ) of the rug be ?

OpenStudy (campbell_st):

let the width of the room be x, then the length is x + 2 so then the area of 168 is \[x \times (x + 2) = 168\] or \[x^2 + 2x - 168 = 0\] solve for x.

OpenStudy (campbell_st):

this will allow you to get the dimensions of the room. then you can get the dimensions of the rug.